# FLRW embeddings in $\mathbb{R}^{n+2}$, differential geometry and conformal photon propagator

### v1

@@ C.3 FLRW spaces with $k=+1$
For the $k=+1$ case, as its defining function resembles that of the k=1k=-1 case, one expects to draw similar conclusions. This is only partially the case. The computations parallels completely those of the k=1k=-1 case, which we skip here, focusing instead on the differences. Very similarly to the k=1k=-1 case we have one additional isometry $y_{0}\partial_{n+1}-y_{n+1}\partial_{0}$ iff $\overset{\,{}_{\mbox{\Large.}}}{a}=0$, that is for $a=1$ which is the $k=+1$ Einstein space. Much like the k=1k=-1 scenario, there exist three specific instances that result in the occurrence of $n$ extra isometries. First, if $a(t)$ fulfills $\tan(t)\overset{\,{}_{\mbox{\Large.}}}{a}+a=0$, that is for $a(t)=1/\sin(t)=\csc(t)$, corresponding to a dS space [Eq. (68)]. Second, if $a(t)$ fulfills $\overset{\,{}_{\mbox{\Large.}}}{a}-\tan(t)a=\overset{\,{}_{\mbox{\Large.}}}{a}+a/\tan(t-\pi/2)=0$ which is yet another instance of a de Sitter space. [^3] Otherwise a direct integration yields $a(t)=\cos(t)$ leading to the metric $\eta_{\scriptscriptstyle{H^{2}}}=\cos^{2}(t)[dt\otimes dt-g_{\mathbb{S}^{n-1}}]=d\tau\otimes d\tau-(1-\tau^{2})g_{\mathbb{S}^{n-1}}$ over de Sitter space. Third, and finally, the equivalent to Eq. (72) here reads as
@@ Appendix D Proofs of various properties
$=\mathrm{sgn}(\eta)(-1)^{n(a+1)+1}\{*_{f}d_{f}m_{f}{}^{*}(i_{\scriptscriptstyle{\,}\!F}i_{\scriptscriptstyle{\,}\!D}*_{\eta}\alpha)+*_{f}m_{f}{}^{*}[(i_{\scriptscriptstyle{\,}\!F}(\mathscr{L}_{\scriptscriptstyle{\!}D}-1)-i_{\scriptscriptstyle{\,}\!D}\mathscr{L}_{\scriptscriptstyle{\!}F})*_{\eta}\alpha]\}$ (73)
@@
$=-m_{f}{}^{*}[\{(\mathscr{L}_{\scriptscriptstyle{\!}D}-2a+n+1)i_{\scriptscriptstyle{\,}\!F}-\mathbf{F}^{2}(\mathscr{L}_{\scriptscriptstyle{\!}D}-2a+n+1)i_{\scriptscriptstyle{\,}\!D}+(\mathscr{S}^{df}+\square_{\eta}f)i_{\scriptscriptstyle{\,}\!D}\}\alpha]$
@@
with a second set of additional terms coming from $(d\mathbf{F}^{2})\wedge\beta=j^{d\mathbf{F}^{2}}\beta$ and from the commutator $d,(\mathscr{S}^{df}+\square_{\eta}f)]i_{\scriptscriptstyle{\,}\!D}\alpha$. ∎
@@
Finally on $X_{f}$, since one has $m_{f}{}^{*}\square_{\eta}\hat{\beta}=\square_{\eta}\hat{\beta}-m_{f}{}^{l}\square_{\eta}\hat{\beta}$, this leads to
@@
$\mathrm{Rm}_{f}=\tfrac{1}{2}[\ell_{n+1}\owedge\ell_{n+1}-\ell_{n}\owedge\ell_{n}]=\tfrac{1}{2}(\ell_{n+1}-\ell_{n})\owedge(\ell_{n+1}+\ell_{n}),$ (75)
@@
and, similarly $g_{bd}T_{ac}j^{a}i^{b}j^{c}i^{d}={}^{\scriptscriptstyle{D}}T$. The third term gives

### v2

@@ C.3 FLRW spaces with $k=+1$
For the $k=+1$ case, as its defining function resembles that of the k=1k=-1 case, one expects to draw similar conclusions. This is only partially the case. The computations parallels completely those of the k=1k=-1 case, which we skip here, focusing instead on the differences. Very similarly to the k=1k=-1 case we have one additional isometry $y_{0}\partial_{n+1}-y_{n+1}\partial_{0}$ if $\overset{\,{}_{\mbox{\Large.}}}{a}=0$, that is $a=1$ and the space is the $k=+1$ Einstein space. Much like the k=1k=-1 scenario, there exist three specific instances that result in the occurrence of $n$ extra isometries. First, if $a(t)$ fulfills $\tan(t)\overset{\,{}_{\mbox{\Large.}}}{a}+a=0$, that is for $a(t)=1/\sin(t)=\csc(t)$, corresponding to a dS space [Eq. (68)]. Second, if $a(t)$ fulfills $\overset{\,{}_{\mbox{\Large.}}}{a}-\tan(t)a=\overset{\,{}_{\mbox{\Large.}}}{a}+a/\tan(t-\pi/2)=0$ which is yet another instance of a de Sitter space. [^3] Otherwise a direct integration yields $a(t)=\cos(t)$ leading to the metric $\eta_{\scriptscriptstyle{H^{2}}}=\cos^{2}(t)[dt\otimes dt-g_{\mathbb{S}^{n-1}}]=d\tau\otimes d\tau-(1-\tau^{2})g_{\mathbb{S}^{n-1}}$ over de Sitter space. Third, and finally, the equivalent to Eq. (72) here reads as
@@ Appendix D Proofs of various properties
$=\mathrm{sgn}(\eta)(-1)^{n(a+1)+1}\{*_{f}d_{f}m_{f}{}^{*}(i_{\scriptscriptstyle{\,}\!F}i_{\scriptscriptstyle{\,}\!D}*_{\eta}\alpha)+*_{f}m_{f}{}^{*}[(i_{\scriptscriptstyle{\,}\!F}(\mathscr{L}_{\scriptscriptstyle{\!}D}-1)-i_{\scriptscriptstyle{\,}\!D}\mathscr{L}_{\scriptscriptstyle{\!}F})*_{\eta}\alpha]\},$ (73)
@@
$=-m_{f}{}^{*}[\{(\mathscr{L}_{\scriptscriptstyle{\!}D}-2a+n+1)i_{\scriptscriptstyle{\,}\!F}-\mathbf{F}^{2}(\mathscr{L}_{\scriptscriptstyle{\!}D}-2a+n+1)i_{\scriptscriptstyle{\,}\!D}+(\mathscr{S}^{df}+\square_{\eta}f)i_{\scriptscriptstyle{\,}\!D}\}\alpha],$
@@
and a second set of additional terms coming from $(d\mathbf{F}^{2})\wedge\beta=j^{d\mathbf{F}^{2}}\beta$ and from the commutator $d,(\mathscr{S}^{df}+\square_{\eta}f)]i_{\scriptscriptstyle{\,}\!D}\alpha$. ∎
@@
Finally, on $X_{f}$, since one has $m_{f}{}^{*}\square_{\eta}\hat{\beta}=\square_{\eta}\hat{\beta}-m_{f}{}^{l}\square_{\eta}\hat{\beta}$, this leads to
@@
$\mathrm{Rm}_{f}=\tfrac{1}{2}(\ell_{n+1}\owedge\ell_{n+1}-\ell_{n}\owedge\ell_{n})=\tfrac{1}{2}(\ell_{n+1}-\ell_{n})\owedge(\ell_{n+1}+\ell_{n}),$ (75)
@@
and, similarly, $g_{bd}T_{ac}j^{a}i^{b}j^{c}i^{d}={}^{\scriptscriptstyle{D}}T$. The third term gives